OpenStudy (anonymous):

sin3x cos2x- cos3x sin 2x Simplify.

OpenStudy (whpalmer4):

Remember your trig identities, specifically \[\sin(u\pm v) = \sin u \cos v \pm \cos u \sin v\] Can you match that up with this problem?

OpenStudy (luigi0210):

Here is a hint, it has an s and a minus in it :P